What thickness, based on deflection, would be chosen for an 18 long simple span beam with Grade 60 steel reinforcing? 1.

this is structural mechanic concrete
Answer #1

Length of span = 18 ft = 18 * 12 = 216 in

Simply supported beam

Grade 60 steel reinforcement

Based on deflection, thickness of beam, h = L/16

h = 216/16 = 13.5 in